that's all they give

isin excess and how many grams of this reactant wil remain after the reaction is complete? O Caco O Ha
Answer #1

CaCO3 + 2HCl -----> CaCl2 + H2O +CO2

1 mole of CaCO3 reacts with 2 moles of HCl and produces 1 mole of CaCl2

Moles of CaCO3 = Mass/Molarmass = 29/100 = 0.29

Moles of HCl = Mass/Molarmass = 13/36.46 = 0.3565

CaCO3 : HCl : CaCl2 = 1 : 2 : 1

[0.29 moles of CaCO3 reacts with 0.29*2 = 0.58 moles of HCl ---> but that much moles of HCl not present ...so limiting rectant is HCl ]

So 0.3565 moles of HCl reacts with 0.3565/2 = 0.17825 moles of CaCO3 and produces 0.17825 moles of CaCl2

Mass of CaCl2 = MW *moles = 110.98*0.17825 = 19.78 gms

Excezz reactant = CaCO3

Excess amount = [0.29-0.17825]*36.46 = 4.0744 gms